a man is 5 times as old as his daughter. the man's son is 2 years older than his sister. in 10 years the sum of their three ages is 81. what are each of their ages?
Answer by lwsshak3(11628)

a man is 5 times as old as his daughter. the man's son is 2 years older than his sister. in 10 years the sum of their three ages is 81. what are each of their ages?
***
at present:
let x=daughter's age
5x=fathers age
x+2=son's age
..
10 yrs hence:
x+10=daughter's age
5x+10=fathers age
x+12=son's age
x+10+5x+10+x+12=81
7x+32=81
7x=49
x=7
5x=35
x+2=9
daughter's age=7
father's age=35
son's age=9